Find the Fourier Sine and Cosine transformations of the following function f (x) =2 if x< 0<a and f(x)=0 if x>a


1
Expert's answer
2021-08-10T06:08:35-0400

Solution,

For the Fourier sine transform,

By definition;

"f_s[{f(x)}]=\hat f_s(w)=\sqrt{\frac2π}\int_0^\infin f(x)sin(wx)dx"

Hence;

"\hat f_s(w)=\sqrt{\frac2π}[\int_0^a2sin(wx)dx+0]"

"\hat f_s(w)=\sqrt\frac2π[\frac{-2cos(wx)}{w}|_0^a]"

"\hat f_s(w)=-\frac2w\sqrt\frac2π(cos(wa)-cos0)"

"\hat f_s(w)=-\frac2w\sqrt\frac2π(cos(wa)-1)"

For cosine Fourier transform,

By definition;

"\hat f_c(w)=\sqrt\frac2π\int_0^\infin f(x)cos(wx)dx"

"\hat f_c(w)=\sqrt\frac2π[\int_0^a2cos(wx)dx+0]"

"\hat f_c(w)=\sqrt\frac2π(\frac{2sin(wx)}{w}|_0^a)"

"\hat f_c(w)=\frac2w\sqrt\frac2π(sin(wa))"